# Embodied Agents Meet Personalization: Exploring Memory Utilization for Personalized Assistance

### v1

@@ B.2 LLM-Powered Embodied Agent Architecture
Following Szot et al. [45], Puig et al. [38], Chang et al. [7], we adopt a two-layer hierarchical control architecture for our LLM-powered embodied agent. We utilize the LLM as a high-level policy planner that selects appropriate skills from the predefined skill library. The selected skill then provides control signals to the simulator. For memory systems, we implement a textual scene-graph as our semantic memory alongside an episodic memory.
@@
For semantic memory, we implement a scene-graph style hierarchical representation, which has demonstrated effectiveness for planning problems [1, 39, 16]. Following Chang et al. [7], we utilize a multi-edge directed graph with three distinct levels to represent environmental entities. The top level contains a single root node representing the house environment, the second level comprises room nodes, and the third level encompasses furniture, objects, and agents. Each node stores the corresponding entity’s 3D location and relevant state information. This graph structure is initialized and continuously updated with ground-truth information from the simulator at each state $s_{t}$. In our system, this structured semantic memory provides the LLM planner with an interpretable representation of the environment, which can be flexibly queried and reasoned over through natural language descriptions.
@@
Our episodic memory is configured to store the ReAct-style formatting that guides the LLMs’ reasoning process [56]. This memory structure captures both the user’s instruction and the complete sequence of <Thought, Action, Observation> triplets generated during task execution. The episodic memory is accessed at the beginning of each task by retrieval and updated upon task completion, enabling the agent to recall previous interactions.
@@
For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer [40] and use the current task instruction as the query. To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
LLM Setup.
@@
B.3 Computing resources
Our experiments primarily utilized commercial API services rather than local computing resources. We used OpenAI’s Chat API for accessing GPT-4o, Claude models through Anthropic’s API, and OpenRouter’s API service for accessing Llama-3.1 [33], Qwen-2.5 [54]. For running simulation environment we used 8 NVIDIA GeForce RTX 3090 GPUs. For our implementation and evaluation, we use Huggingface library2, vLLM library. Both libraries are licensed under Apache License, Version 2.0. And we used langchain library, under MIT License. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
Appendix C Details of Memento
C.1 Personalized Knowledge
@@
C.2 PartNR Dataset
@@
PartNR is designed to evaluate planning and reasoning capabilities in embodied tasks and is recognized for its comprehensive collection of natural language instructions in household environments. The benchmark includes four primary task types: (1) constraint-free basic rearrangement tasks, (2) spatial tasks requiring reasoning about object positions, (3) temporal tasks with sequential dependencies, and (4) heterogeneous tasks involving actions that can only be performed by human agents. We selected PartNR episodes specifically for their complexity beyond simple pick-and-place operations. The rich linguistic structure and diverse task requirements make PartNR particularly suitable for evaluating user patterns personalization, enabling us to create scenarios that effectively test an agent’s ability to adapt to user-specific communication patterns.
C.3 Dataset Statistics
@@
C.4 Data Generation Process
To create stage-specific tasks, we leveraged GPT-4o and systematic process to incorporate personalized knowledge into existing task structures.
Captioning process.
Since we employ LLMs as high-level planners for embodied agent, we generated natural language descriptions of the objects at the scene using GPT-4o, to enable agents to reason over object descriptions without relying on direct visual perception. We collected object models from the OVMM dataset [57], and used GPT-4o to generate natural language descriptions from rendered object images. Especially, the Google Scanned dataset [14], included within OVMM, provides object identities in file names. We leveraged this additional information alongside the images to produce more realistic descriptions. Through this process, we generated 1,920 object descriptions with 66 categories of objects. The prompts used to generate the object descriptions are provided in Appendix E.
Preprocessing scenes.
To collect the suitable episodes for our purposes, we preprocessed and filtered episodes. First, we only gathered non-heterogeneous episodes, which should be executed by human agents. Second, we filtered out tasks where the target object was not uniquely specified (e.g., “Bring one apple to the kitchen table”), as our task setting requires identification and reference of a specific object based on personalized knowledge that distinguishes it from other similar objects. Third, in cases where captions for target handles were unavailable, we substituted alternative objects. If no objects from the same category were available, we excluded the entire episode from our dataset.
Distractor sampling.
To sample distractor objects for episodes focusing on object semantics, we utilized PartNR’s dataset generation methods. This approach allowed us to systematically select objects located adjacent to target objects on the same receptacles or floor surfaces, requiring agents to differentiate between objects of the same category.
Details of task instruction generation.
For instruction generation, we prompted GPT-4o to generate personalized knowledge tasks based on the knowledge categories defined in Section C.1. For tasks involving object semantics, we provided the captions of the target object pairs along with instructions to guide GPT-4o in generating natural object semantics descriptions. (e.g., instruction: "Bring the cup on the kitchen table", object description: "a white mug with fancy handle", personalized knowledge: "The mug is my coffee mug.") We first instructed the model to create the most plausible subcategory of object semantics, which was then used to generate personalized knowledge specific to the objects. In these object semantics cases, the instruction comprised three components: command instruction (e.g., "Bring the cup on the kitchen table"), additional information (e.g., "The cup is a white mug with fancy handle."), and personalized knowledge (e.g., "That mug is my coffee mug"). For user patterns tasks, we provided instructions to GPT-4o and allowed the model to infer the most plausible user patterns corresponding to a sequence of actions. For the memory acquisition stage instruction, we also concatenated the command instruction (e.g., "Bring the cup and dish from the kitchen table to living room") with personalized knowledge (e.g., "That’s my dinner setup"). Detailed information about the prompts we used is provided in Appendix E.
Quality control.
After generating episodes for Memento, we implemented a two-stage filtering process for quality control. First, we applied heuristic filtering to eliminate episodes with potentially confusing distractor objects in the same scene (e.g., episodes using identical object handles or requiring similar personalized knowledge, which could create ambiguity). Second, we tested each episode with GPT-4o and excluded any episode where GPT-4o failed five consecutive attempts, indicating potential issues with task feasibility or clarity. (Figure 7) This quality control process resulted in filtering out 31 episodes (13.4% of the total dataset) that exhibited consistently poor performance. Specifically, we identified: 31 episodes with zero success rate, indicating complete task failure. The remaining 201 episodes showed strong performance metrics, with 95% confidence intervals of [0.755, 0.844] for success rate and [0.823, 0.893] for completion rate. The high correlation (r = 0.908) between success rate and completion rate indicates consistent performance across different evaluation metrics. These high-quality episodes served as the foundation for generating golden trajectories, which were subsequently used in our memory quality analysis and discussion experiments. By establishing GPT-4o’s successful task executions as golden trajectories, we ensure that our evaluation framework is based on demonstrably achievable performance standards, providing a reliable benchmark for assessing memory quality and discussion effectiveness.
Appendix D Additional Experiments & Analysis
D.1 Knowledge Type-based Analysis.
Additional experiment results show the performance of small models—Llama3.1-8b and Qwen2.5-7b—on knowledge type-based analysis (Figure 8). Similar to other frontier models, they struggled with utilizing episodic memory on tasks requiring recognition of user patterns. For tasks requiring object semantics during the memory acquisition stage, we found that these small models struggle to understand the user’s intent to differentiate objects with DescribeObjectTool and rely primarily on commonsense reasoning, instead of describing the objects.
Figure 8: Memory acquisition stage and memory utilization stage all models.
D.2 Knowledge Type-based Analysis on Dual-Memory Tasks
Figure 9 shows the success rate of each model on dual-memory tasks, compared with the discrete success/failure outcomes when executing each corresponding episode individually during the memory acquisition stage. We can observe that almost all models tend to struggle with utilizing user patterns knowledge from different memory sources, with smaller models (Llama3.1-8b and Qwen2.5-7b) demonstrating particularly pronounced difficulties.
Figure 9: Personalized knowledge type-based analysis on joint-memory tasks, comparing with memory acquisition stage’s corresponding episodes.
D.3 Memory Quality Analysis

### v2

@@ C.1 LLM-Powered Embodied Agent Architecture
Following Szot et al. (2021); Puig et al. (2023); Chang et al. (2024), we adopt a two-layer hierarchical control architecture for our LLM-powered embodied agent. We utilize the LLM as a high-level policy planner that selects appropriate skills from the predefined skill library. The selected skill then provides control signals to the simulator. For memory systems, we implement a textual scene-graph as our semantic memory alongside an episodic memory.
@@
For semantic memory, we implement a scene-graph style hierarchical representation, which has demonstrated effectiveness for planning problems (Agia et al., 2022; Rana et al., 2023; Gu et al., 2024). Following Chang et al. (2024), we utilize a multi-edge directed graph with three distinct levels to represent environmental entities. The top level contains a single root node representing the house environment, the second level comprises room nodes, and the third level encompasses furniture, objects, and agents. Each node stores the corresponding entity’s 3D location and relevant state information. This graph structure is initialized and continuously updated with ground-truth information from the simulator at each state $s_{t}$. In our system, this structured semantic memory provides the LLM planner with an interpretable representation of the environment, which can be flexibly queried and reasoned over through natural language descriptions.
@@
Our episodic memory is configured to store the ReAct-style formatting that guides the LLMs’ reasoning process (Yao et al., 2023). This memory structure captures both the user’s instruction and the complete sequence of <Thought, Action, Observation> triplets generated during task execution. The episodic memory is accessed at the beginning of each task by retrieval and updated upon task completion, enabling the agent to recall previous interactions.
@@
For retrieval, we encode instructions and memory entries using the all-mpnet-base-v2 sentence transformer (Reimers & Gurevych, 2019) and use the current task instruction as the query. To avoid ambiguous or irrelevant memories, we retrieve candidate memories only from the history within the same scene as the current task.
LLM inference setup.
@@
C.2 Personalized Knowledge Details
@@
C.3 Details of Memento
C.3.1 Foundation Dataset
@@
We adopt the test set of PartNR (Chang et al., 2024) as our foundation object rearrangement task dataset. PartNR is designed to evaluate planning and reasoning capabilities in embodied tasks and includes four primary task types: (1) constraint-free basic rearrangement, (2) spatial tasks requiring reasoning about object positions, (3) temporal tasks with sequential dependencies, and (4) heterogeneous tasks involving human-only actions. We select PartNR episodes for their complexity beyond simple pick-and-place operations, filtering out tasks irrelevant to personalized assistance scenarios. The rich linguistic structure and diverse task requirements make PartNR particularly suitable for evaluating personalized user patterns, enabling us to create scenarios that effectively test agents’ ability to adapt to user-specific behavioral routines.
C.3.2 Dataset Generation Process Details
We leverage GPT-4o and a systematic process to incorporate personalized knowledge into existing task structures, creating tasks tailored for each evaluation stage.
Captioning process.
Since we employ LLMs as high-level policy planners for embodied agents, we generate natural language descriptions of objects in the scene using GPT-4o, enabling agents to reason over object properties without relying on direct visual perception. We collect object models from the OVMM dataset (Yenamandra et al., 2023) and use GPT-4o to generate descriptions from rendered object images. In particular, the Google Scanned dataset (Downs et al., 2022) included within OVMM provides object identities in file names, which we leverage alongside the images to produce more realistic and accurate descriptions. Through this process, we generate 1,920 object descriptions spanning 66 object categories. The prompts used for generating object descriptions are provided in Appendix F.
Preprocessing scenes.
To collect suitable episodes for personalized assistance tasks, we preprocessed and filtered episodes using three criteria. We only gathered non-heterogeneous episodes that can be executed by robotic agents, excluding tasks that require human-only actions. We filtered out tasks where target objects were not uniquely specified (e.g., "Bring one apple to the kitchen table"), since our task setting requires identification of specific objects based on personalized knowledge that distinguishes them from similar objects. When captions for target objects were unavailable, we substituted alternative objects from the same category, or excluded the entire episode if no suitable alternatives existed.
Distractor sampling.
To sample distractor objects for episodes focusing on object semantics, we utilized PartNR’s dataset generation methods. This approach systematically selects objects located adjacent to target objects on the same receptacles or floor surfaces, requiring agents to differentiate between objects of the same category.
Task instruction generation process.
We prompt GPT-4o to generate personalized knowledge tasks based on the knowledge categories defined in Section C.2.
For object semantics tasks, we provide target object captions and guide GPT-4o to generate natural semantic descriptions. We first instruct the model to create the most plausible subcategory of object semantics, then generate corresponding personalized knowledge. The resulting instruction comprises three components: the command instruction, additional object information, and personalized knowledge. For example: "Bring the cup on the kitchen table" (command) + "The cup is a white mug with fancy handle" (description) + "That mug is my coffee mug" (personalized knowledge).
For user pattern tasks, we provide base instructions and allow GPT-4o to infer plausible user patterns corresponding to action sequences. Memory acquisition stage instructions concatenate the command with personalized knowledge (e.g., "Bring the cup and dish from the kitchen table to living room. That’s my dinner setup"). Detailed prompts are provided in Appendix F.
Quality control.
To ensure dataset quality, we implement a two-stage filtering process after generating episodes for Memento. We first apply rule-based heuristic filters guided by six explicit criteria to identify problematic episodes:
  • Semantic misalignment – Instructions inconsistent with stored memory (e.g., memory states "my favorite cup" but instruction refers to "my cup")
  • Triviality – Instructions solvable without memory access (e.g., "Prepare my coffee time cup setup on the kitchen table" when setup details are explicitly provided)
  • Redundancy – Instructions repeating knowledge from previous episodes in the same scene (e.g., "my usual morning cup" when an earlier episode already established "the cup I usually use in the morning")
  • Environment inconsistency – Instructions conflicting with actual object properties (e.g., describing target cup as red when it’s actually blue, but a distractor cup is red)
  • Instruction inconsistency – Semantic inconsistency between acquisition and utilization stages (e.g., changing from "pick from A and place at B" to only "place at B")
  • Temporal irrelevance – Instructions with expired temporal context (e.g., referencing "today’s temperature" when the memory is no longer current)
Following this rule-based filtering, we test each remaining episode with GPT-4o, excluding episodes where the model fails five consecutive attempts as this indicates potential feasibility or clarity issues. This comprehensive quality control process filters out 31 episodes (13.4%) showing consistently poor performance, all with zero success rates indicating complete task failure. The remaining 201 episodes demonstrate strong performance with 95% confidence intervals of [0.755, 0.844] for success rate and [0.823, 0.893] for completion rate. The high correlation (r = 0.908) between these metrics confirms consistent performance across evaluation measures.
C.3.3 Inter-Rater Reliability Analysis
To validate the consistency of our heuristic filtering process, we conduct inter-rater reliability analysis using Krippendorff’s alpha (Krippendorff, 2011). Multiple annotators independently evaluate a held-out subset of instructions using our six filtering criteria. Since Environment inconsistency and Temporal irrelevance are strictly controlled during instruction generation, they achieve near-perfect agreement and we exclude them from reliability analysis. We compute Krippendorff’s alpha for the remaining four categories, obtaining values of 0.75 for Semantic misalignment, 0.87 for Instruction inconsistency, 0.58 for Triviality, and 0.57 for Redundancy. The average alpha of 0.69 across these criteria indicates moderate agreement, supporting the reliability of our filtering process for maintaining dataset quality.
C.3.4 Dataset Statistics
@@
C.4 Evaluation Method
We adopt the official evaluation protocol from the PartNR benchmark (Chang et al., 2024), which provides a python-based framework for assessing multi-step rearrangement tasks. We use this framework without modification. The evaluator analyzes simulator states using three components: (1) propositions that define object relationships to satisfy (e.g., is_on_top([spoon_1], [table_1])), (2) dependencies that define temporal conditions for multi-step instructions (after_satisfied, after_unsatisfied), and (3) constraints that enforce execution requirements (e.g., ordering, object consistency across steps). We rely on this system to evaluate tasks involving ambiguous references and sequential dependencies. The evaluation produces a percent-complete score and binary success indicator.
C.5 Computing resources
Our experiments primarily utilized commercial API services rather than local computing resources. We used OpenAI’s Chat API for accessing GPT-4o, Claude models through Anthropic’s API, and OpenRouter’s API service for accessing Llama-3.1 (Meta, 2024), Qwen-2.5 (Yang et al., 2024). For running simulation environment we used 8 NVIDIA GeForce RTX 3090 GPUs. For our implementation and evaluation, we use Huggingface library2, vLLM library. Both libraries are licensed under Apache License, Version 2.0. And we used langchain library, under MIT License. We have confirmed that all of the artifacts used in this paper are available for non-commercial scientific use.
Appendix D Experiment & Analysis Details
In this section, we provide a detailed description of the experimental setup, analysis results, and the memory module, complementing the explanations provided in the main text.
D.1 User Profile Memory
Our work introduces a user profile memory module and offers initial evidence that, when combined with episodic memory, structured personalized knowledge can enhance performance on personalized assistance tasks.
Knowledge graph design. We define the user profile memory as a hierarchical knowledge graph that organizes personalized knowledge into three levels—User, Knowledge, and Elements—to provide a structured representation that is both scalable across users and resistant to redundancy from overlapping entities. The hierarchical partitioning enables the system to scale to many users while preserving efficiency, since overlapping objects can be represented in a canonical form and connected via edges instead of being redundantly replicated. To support this representation, each nodes are connected through two types of edges: hierarchical edges, which capture inclusion relations across and within levels, and temporal edges, which encode the sequential structure of user patterns. Together, these edge types allow the graph to comprehensively represent the personalized knowledge that users may require an embodied agent to leverage in real-world interactions.
Input: User instruction II, Knowledge graph $G$
Output: Updated knowledge graph $G^{\prime}$
// Step 1: Extract candidate knowledges from instruction
$\mathcal{K}\leftarrow ExtractKnowledge(I)$ // $\mathcal{K}=\{(k_{1},t_{1}),(k_{2},t_{2}),\dots\}$
$C\leftarrow\varnothing$
1ex
foreach $(k,t_{k})\in\mathcal{K}$ do
 // Step 2: Extract elements (patterns/objects/locations) tied to this knowledge
 $\mathcal{E}\leftarrow ExtractElements(I,k,t_{k})$;
 // Step 3: Select subgraph by type
 $G_{t_{k}}\leftarrow Subgraph(G,type=t_{k})$;
 // Step 4: Retrieve candidates for knowledge and elements
 $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},k,t_{k})$;
 foreach $(e,t_{e})\in\mathcal{E}$ do
    $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},e,t_{e})$;
// Step 5: Use LLM to decide add or update
$op\leftarrow LLMResolve(I,C)$;
if $op=\mathtt{update}$ then
 return $G^{\prime}\leftarrow UpdateKnowledgeNode(G,Expand(C),I)$;
else if $op=\mathtt{add}$ then
 return $G^{\prime}\leftarrow AddKnowledgeNode(G,Expand(C),I)$;
Algorithm 1 Update Knowledge Graph
Knowledge graph update procedure. While the hierarchical design specifies how knowledge is organized, the graph must also evolve as users provide new information or refine existing entries. Algorithm 1 presents the procedure for updating the knowledge graph. Given a user instruction $I$ and the current graph $G$, the procedure begins by extracting candidate knowledge $\mathcal{K}$ and associated elements $\mathcal{E}$. For each candidate knowledge $k$, the relevant subgraph $G_{t_{k}}$ is filtered, and similarity search is applied to collect matching nodes $C$. These candidates are then expanded by collecting all connected edges and their descendant nodes, and the resulting subgraphs are passed to an LLM. The LLM determines whether the instruction corresponds to an addition or an update and performs the corresponding operation, yielding the updated graph $G$’.
Input: User instruction II, User profile memory graph $G$
Output: Retrieved natural language descriptions $R$
// Step 1: Parse instruction
$\mathcal{K}\leftarrow ExtractKnowledge(I)$;
$\mathcal{C,R}\leftarrow\varnothing$
// Step 2: Filter subgraph and run similarity search
foreach $(k,t_{k})\in\mathcal{K}$ do
 $G_{t}\leftarrow Subgraph(G,type=t)$;
 $C\leftarrow C\cup SimilaritySearch(G_{t_{k}},k,t_{k})$;
$C\leftarrow RemoveDuplicates(C)$;
// Step 3: Reformulate results with LLM
foreach $(c,t_{c})\in\mathcal{C}$ do
 $R\leftarrow R\cup Reformulate(Expand(c))$;
return $R$
Algorithm 2 RetrievalProcedure
Retrieval procedure. To leverage personalized knowledge during task execution, the agent retrieves information from the user profile memory using a hybrid retrieval pipeline that combines embedding-based search with LLM reasoning. This procedure ensures that the retrieved knowledge is both semantically aligned with user instructions and interpretable in natural language.
Knowledge graph schema. We summarize the node and edge types in Tables 7 and 8. Complete examples of user profile memory instances are shown in Figures 12 and 13.
Table 7: Edge schema of the user profile memory, including relation types and JSON-style examples.
Edge Type Source Relation Target Example
Hierarchical user refers_to knowledge { "source": "u1", "target": "k1", "type": "Hierarchical", "relation": "refers_to" }
Hierarchical knowledge entails pattern { "source": "k1", "target": "p1", "type": "Hierarchical", "relation": "entails" }
Hierarchical pattern target object { "source": "p1", "target": "o1", "type": "Hierarchical", "relation": "target" }
Hierarchical pattern target location { "source": "p1", "target": "l1", "type": "Hierarchical", "relation": "target" }
Hierarchical knowledge composed_of object { "source": "k1", "target": "o2", "type": "Hierarchical", "relation": "composed_of" }
Temporal pattern before pattern { "source": "p1", "target": "p2", "type": "Temporal", "relation": "before" }
Table 8: Node schema of the user profile memory.
Node Type Attributes Example
User id, type, name {u1, "user", "james"}
Knowledge id, type, subtype, alias, description {k_obj1, "knowledge", object_semantics, "my favorite cup", "a yellow cup with a wooden handle"}
Pattern id, type, name, args {p1, "pattern", "place", [cup, on, livingroom_table]}
Object id, type, name, granularity {o2, "object", "red dog image cup", instance}
Location id, type, name, expression {l1, "location", livingroom_table, "on the table"}
D.2 Success and Error Case Analysis
Tasks require object semantics knowledge.
Success cases demonstrate that agents can effectively reference personalized object attributes from episodic memory, correctly identifying and applying the specific information needed for task completion. (A.1) However, we observed distinct error patterns when agents needed to utilize this type of knowledge: the most prominent being commonsense reasoning errors (B.1), where agents incorrectly relied on generic object descriptions instead of personalized cues; missed personalization cues (B.2), where agents failed to recognize the need to access personalized knowledge; and memory recall failures (B.3), where agents were unable to locate relevant information despite its presence in the provided context.
Tasks require user patterns.
D.3 Memory Design Experiment Details
We describe here the experimental setup for Section 6. The following details clarify the episode sampling, memory construction, and retrieval configurations used in this experiment.
D.3.1 Details of Memory Format Simplification Experiment
We sampled 30 episodes by selecting 10 episodes from each of three scenes, with careful balancing across task types and difficulty levels informed by our preliminary experimental results. To minimize the influence of erroneous trajectories, we provided GPT-4o-generated gold memory as the given episodic memory, thereby reducing the impact of noisy interaction histories. We also used GPT-4o to generate summaries of the action trajectories, including user instructions, to facilitate compact memory representations. For each evaluation, one exemplar shot was given to the agent, and top-$k$ memory entries were retrieved based on the current query, where $k=10$ for frontier models (GPT-4o and Qwen-2.5-72B) and $k=7$ for smaller models (LLaMA-3.1-8B and Qwen-2.5-7B), considering context length limitations. The prompts used for these experiments are provided in Appendix F.3.
D.3.2 Details of User Profile Memory Experiment
We conduct experiments by incorporating user profile memory in addition to episodic memory. The experiments are conducted with the same models as in Section 6.1 (GPT-4o, Qwen-2.5-72B-Instruct, Llama-3.1-8B-Instruct, and Qwen-2.5-7B-Instruct). These models are also employed both to construct the user profile memory and to carry out the retrieving procedure, which combines embedding-based search with LLM reasoning: the LLM parses user instructions and reformulates retrieved graph content into natural language descriptions. User profile memory entries are retrieved under a $k=5$ setup without gold memory, while episodic memory is supplied as each agent’s trajectory under a $k=5$ setup with gold memory. We evaluate performance on both the single and joint memory utilization stages, and the prompts for memory construction and retrieval are provided in Appendix F.
D.4 Example Trajectory and Evaluation
To provide readers with a concrete understanding of our evaluation framework and typical agent performance patterns, we present detailed examples of both successful (Tables 161718) and failure cases (Tables 192021). These examples include the complete instruction metadata, agent execution trajectories, and evaluation outcomes, offering insights into when and why agents succeed or struggle with personalized knowledge utilization.
Appendix E Extended Analysis: Memory Components and Real-World Scenarios
To provide comprehensive insights into memory utilization for personalized assistance, we conduct extended analysis examining multiple factors affecting agent performance. This section covers three primary areas: memory quality effects on agent performance, retrieval strategy effectiveness for episodic memory, and agent behavior in complex real-world scenarios including 3-joint memory tasks and ambiguous instructions.
E.1 Human Baseline
Experiment setup.
To establish a meaningful comparison point for our agent evaluation and validate that our tasks are fundamentally solvable, we conduct a human baseline study. We recruit four undergraduate participants to complete a curated set of 20 tasks that represent the challenges our embodied agents face in memory utilization for personalized assistance.
To ensure comprehensive coverage of task complexity, we categorize task difficulty based on empirical observations of model behavior across our agent evaluations:
  • Easy: Llama-8B successfully completed the task using its own memory
  • Medium: Llama-8B failed with its own memory but succeeded when given GPT-4o’s answer memory
  • Hard: Llama-8B failed even when given GPT-4o’s answer memory
  • Super hard: GPT-4o failed with its own memory
  • Random joint-memory tasks
We implement a text-based interactive interface that mirrors the agent environment while accommodating human interaction patterns. Participants issue step-by-step commands and receive observational feedback after each action, similar to how our agents operate. Each participant receives a 10-minute tutorial on the interface and completes tasks in the same environment as the agents. Crucially, participants have access to the same top-5 retrieved memories that agents receive, allowing direct comparison of memory utilization capabilities.
Results.
The human baseline provides both validation and perspective on our evaluation framework. All participants successfully completed 100% of the tasks with correct end states, demonstrating that the challenges are fundamentally solvable given appropriate memory access and reasoning capabilities. While humans require more time than LLM-based planners, they exhibit superior abilities in referencing episodic memory, recognizing errors, and adapting their actions accordingly. This performance establishes a meaningful upper-bound reference for agent capabilities and confirms that our tasks capture genuine challenges in memory utilization for personalized assistance.
E.2 Memory Quality and Robustness Analysis
To better understand the factors affecting memory utilization performance, we conduct additional analysis examining two critical aspects of memory-based personalized assistance. First, we analyze how memory quality impacts agent performance by comparing gold memory trajectories with retrieved memory from actual agent interactions. Second, we evaluate agent robustness when episodic memories are corrupted or incomplete through various noise and degradation conditions. These analyses provide insights into the practical limitations and requirements for deploying personalized embodied agents in real-world scenarios where memory quality cannot be guaranteed.